the mom is 32 years old and the daughter is 8 years old im how many years is the mom 3 times as old as the daughter
Answer by graphmatics(170)

Let x be the number until the mom is 3 times as old as the daughter. After x years the daughter is 8+x and the mom is 32+x So 3*(8+x)=32+x. 24+3*x=32+x. 2*x=8. X=4...